# Next-Gen Software Engineering: AI-Assisted Big Models

### v1

@@ Abstract
The effectiveness of model-driven software engineering (MDSE) has been demonstrated in the context of complex software; however, it has not been widely adopted due to the requisite efforts associated with model development and maintenance, as well as the specific modelling competencies required for MDSE. Concurrently, artificial intelligence (AI) methods, particularly machine learning (ML) methods, have demonstrated considerable abilities when applied to the huge code bases accessible on open-source coding platforms. The so-called big code provides the basis for significant advances in empirical software engineering, as well as in the automation of coding processes and improvements in software quality with the use of AI. The objective of this paper is to facilitate a synthesis between these two significant domains of software engineering (SE), namely models and AI in SE. The paper provides an overview of the current status of AI-assisted software engineering. In light of the aforementioned considerations, a vision of AI-assisted Big Models in SE is put forth, with the aim of capitalising on the advantages inherent to both approaches in the context of software development. Finally, the new paradigm of pair modelling in MDSE is proposed.
@@
Software Engineering, Model-Driven Engineering, Artificial Intelligence, Big Models, AI4SE
s
@@ 1 Introduction
Software engineering (SE) is a field of informatics/computer science that addresses the development and analysis of systematic approaches for the design, development, verification & validation and maintenance of software and of software-based systems [^1] In this paper, the term ”software” will be used as a general reference to the collective term for computer programs and related SE artefacts.. It establishes principles and identifies optimal practices for the production and operation of software, with the aim of ensuring the reliability, security, scalability, maintainability and alignment of software with user, business and societal requirements. The objective of software engineering is the production of high-quality software that is cost-effective, delivered in a timely manner and can be readily adapted as requirements evolve.
@@
The term software engineering was coined in 1967 with the very first software crisis Valdez (1988), but still software engineering practices and the resulting software quality have not kept pace with the quality levels required in critical domains or application contexts. Recently, another dramatic story was added to the software horror show: The root cause analysis for the millions of outages caused by the CrowdStrike Falcon sensor CrowdStrike (2024) found that the number of fields in an IPC template type was not validated, a runtime array bounds check was missing, the content validator contained a logic error, template type testing was too limited, and template instances were not tested within the content interpreter. This was unprofessional software development that ignored state of the art of SE.
@@
However, software outages are not the only indication that SE still has no full answers to the challenges posed by the increasing complexity of software-based systems and the diversity of requirements placed on them: Shah et al. (2012) analysed that “[t]here exists a statistically significant medium sized difference between open and closed source projects: the former have a $D$ [(defect density)] that is 4 defects per KLoC lower than the latter. Java projects exhibit a significantly lower $D$ than C projects, 4.1 defects per KLoC on average: In general the Size appears to be negatively correlated to $D$: the larger the project the lower the $D$. In particular, large projects are 10 times less defective than medium ones.”. According to McConnell (2004), the defect density ($D$) of software projects within the industrial sector is estimated to fall within the range of 1 to 25. This suggests that a software program comprising one million lines of code (MLoC) may contain between 1,000 and 25,000 defects. Given the extensive research conducted on the subject of software size in terms of lines of code, the derivation of estimates regarding the approximate number of defects in a software is a relatively straightforward process. To illustrate, Softonic (2023) compares typical software such as operating systems, browsers, and office suites: The Windows 10 operating system has approximately 80 MLoC, Ubuntu 50 MLoC, MacOS X 84 MLoC, Android 12 MLoC, iOS 12 MLoC, the browser Google Chrome 6.7 MLoC, Mozilla Firefox 21 MLoC, or the office suites Microsoft Office 2013 45 MLoC, Apache OpenOffice 19 MLoC, and LibreOffice 10 MLoC. Therefore, the challenges associated with the development of accurate and suitable software remain significant.
@@
Brooks and Bullet (1987) argued that the primary challenge in software development lies in the specification, design, and testing of complex systems, rather than in the labour for representation of the system or testing its fidelity. It emphasized the distinction between essential difficulties (inherent complexity) and accidental difficulties (extraneous challenges), stating that past advances, like high-level programming languages, have only reduced the latter. The paper suggested that addressing accidental difficulties would not lead to major improvements in software development, as essential difficulties remain fundamental. Fraser et al. (2007) critiqued the notion of "accidental difficulties", arguing that these so-called accidents are often the result of negligence or poor practices, not mere happenstance. This paper advocated for a disciplined, science-based, and model-driven approach to software development, similar to traditional engineering disciplines. Building on this, Schieferdecker (2020) highlighted the growing societal dependence on autonomous and intelligent software systems, suggesting that new approaches are needed to ensure not only traditional software quality facets (safety, security, etc.), but also to address socio-technical and socio-political implications, particularly in human-machine collaboration.
@@
In model-driven development (MDD Selic (2003)), also known as model-driven software engineering Schieferdecker (2024), models are original artefacts that are engineered with the intention of facilitating the top-down construction of complex software. Furthermore, models are utilised during runtime to enable the monitoring, verification & validation of software operations, as well as the optimisation of its performance Hailpern and Tarr (2006); France and Rumpe (2007); Weißleder and Lackner (2013).
@@
As has been the case with the comprehensive incorporation of modelling in software engineering, artificial intelligence (AI) methodologies have been employed in software engineering activities from the outset, see for example Barstow (1988). In the context of significant advancements in AI-driven software engineering in 2021, Gartner’s hype cycle for emerging technologies Burke et al. (2021) projected that AI-enhanced software engineering would reach the peak of inflated expectations within a five-to-ten-year timeframe, subsequently entering a phase of productivity stabilization. Since 2023, it is estimated that AI-augmented software engineering will reach the productivity plateau already in two to five years, while in previous iterations of the hype cycle, AI-augmented development and AI-assisted design were identified as emerging technologies on the rise. It was interesting to observe the evolving business expectations and the convergence of trends for AI-supported design and development into a unified technological trend for AI-supported engineering. Moreover, it is encouraging to observe that software engineering is receiving particular attention and prioritisation, which is indicative of the growing importance of this field.
@@
This paper addresses the role of models in software engineering in Section 2, the evolution of big code on coding platforms in Section 3, and the utilisation of AI for coding and software engineering in Section 4. It presents a novel taxonomy of AI for software engineering (AI4SE) to facilitate a deeper understanding of this evolving field. Section 5 examines the practice of contributing software models to coding platforms, which is increasingly being done with software code. In line with the increasing prevalence of large-scale models, it examines recent developments in AI-enhanced modelling and model utilisation within the domain of software engineering. It also presents an updated version of the AI4SE taxonomy, designated as AI4BM, to outline the current status and future prospects of AI applications for Big Models. An outlook concludes the paper.
@@ 2 Models in Software Engineering
According to IEEE (2022) (see Figure 1), the software lifecycle is comprised of distinct phases and activities that can be described as knowledge areas, including requirements, architecture, design, construction, testing and maintenance. There are also knowledge areas that deal with the fundamentals of computer science, mathematics and engineering, as well as cross-cutting activities pertaining to software quality, security, configuration management and engineering management. Furthermore, there are cross-cutting knowledge areas that encompass the processes, operations, professional practice and economics of software engineering, along with a distinct area for software engineering models and methods. This knowledge area encompasses the processes and methodologies associated with modelling, the various types of models, including those pertaining to information, behaviour and structure, and the analysis of models, see also Schieferdecker (2024).
Figure 1: Software Engineering knowledge areas according to SWEBOK IEEE (2022)
Models in SE are the key to dealing with the complexity of software. Models provide the essential abstractions to capture requirements, support design decisions, or to offer comprehensive overviews of software structures and behaviours. They are essential tools in the engineering process for constructing and maintaining software. They are also vital for configuration, monitoring and other runtime support during software operation and management.
The traditional understanding of model-driven software engineering (MDSE) evolved over time:
@@
    In addition to explicit model development and top-down model use, models are also extracted from software executions with the objective of avoiding the costs associated with model development on the one hand and leveraging the benefits of MDSE on the other hand Bagheri and Sullivan (2013).
@@
    The combination of model development of model artefacts with model extraction from code artefacts has resulted in top-down/bottom-up MDSE approaches Steffen et al. (2007)benefiting from both directions of MDSE as shown in Vaupel et al. (2015); Weißleder and Lackner (2013) or Garcia et al. (2023).
@@
    Furthermore, as the concepts underlying programming became increasingly abstract, low-level information and structure models have also become integral to conventional programming. This is exemplified by the use of the data structures often implicitly defined in JSON Lv et al. (2018), XML schemata, or SQL table specifications.
These practical outcomes of model-driven, model-based and/or model-like software developments established a foundation for the utilisation of AI in software engineering, as discussed in Section 5. Prior to this, however, a more detailed view of the current state of practice of big code on open source platforms will be undertaken in Section 3.
@@ 3 Big Code in Software Engineering
The advent of software coding platforms can be traced back to the 1990s, although it was not until the early 2000s that they truly gained prominence, with the launch of Github in April 2008 representing a significant milestone in this regard. At the present time, GitHub is the most widely utilised software coding platform. As stated by Blog (2021), GitHub hosts over 100 million projects and 40 million users. It became not only a significant platform for software engineering collaboration, but also a prominent reference for open-source software mining Kalliamvakou et al. (2014). The study also demonstrated that a considerable number of GitHub repositories are not directly related to software development. This is because GitHub is not solely utilised for coding collaborations; it is also employed for collaborations on websites, editing books or other publications, and is even used as a storage platform. A subsequent manual analysis Kalliamvakou et al. (2016) revealed that over a third of the repositories on GitHub were not software developments.
Nevertheless, the software engineering community can further enhance its comprehension of software engineering processes, methods and tools by gaining insights via the mining of coding platforms. An overview of the GitHub platform as a whole or of a number of repositories can be obtained by utilising tools such as Kibble Apache (2024). In all lines of code of the coding projects on GitHub, approximately 70% are comprised of actual code, with approximately 20% consisting of information structures in formats such as JSON or XML (see Figure 2).
Figure 2: Coding projects breakdown on GitHub by Kibble Apache (2024)
@@
In this context, the concept of "Big Code", analogous to the notion of "Big Data", was introduced in reference to the vast quantities of software code that have accumulated over time Markovtsev and Long (2018); Ortin et al. (2016); Vechev et al. (2016) and that can be analysed and reused to gain deeper insights into software engineering in general, as well as using the code to train ML models.
4 AI in Software Engineering
@@
The use cases for artificial intelligence (AI) in software engineering (SE) are numerous and cover a wide spectrum as discussed already decades ago Barstow (1988) and since then. For example, Feldt et al. (2018) proposes the AI in SE Application Levels (AI-SEAL) taxonomy, which differentiates between the point of applying AI to the software engineering process, the software product or at runtime, the levels of automation in between 1 (“[h]uman considers alternatives, makes and implements decision”) and 10 (“[c]omputer makes and implements decision if it feels it should, and informs human only if it feels this is warranted.”) [^2] Expressing emotion in a computer is currently scientifically problematic because there is no such thing as a computer with the capacity for emotion. A better wording would have been “[c]omputer makes and implements decision if it decides it should, and informs human only if it decides this is warranted.”, and the types of AI along the five tribes differentiation by Domingos (2015). It is noteworthy that the majority of the papers analysed in Feldt et al. (2018) employ AI in the software engineering process. However, despite the numerous facets of the software engineering process (Figure 1), this study does not provide further elaboration. Nevertheless, other publications offer more detailed discussions of AI in SE, including Barenkamp et al. (2020) or Ozkaya (2023). Notwithstanding the aforementioned considerations, a taxonomy for the role of AI in software engineering has yet to emerge.
@@
In light of the recent advancements in machine learning (ML) that have made significant breakthroughs possible, this paper focuses mainly on the latest developments in the application of ML to SE. For ML to be effective, it is essential to utilise the appropriate structures inherent to SE artefacts and has been a topic of considerable debate: Allamanis et al. (2018) presents an overview of the various ways in which source code can be represented, including representational models of tokens, token contexts, program dependency graphs, API calls, abstract syntax trees, object usage, and others. These representational models are used for AI assistance in SE, including the creation of recommender systems, the inference of coding conventions, the detection of anomalies and defects, the analysis of code, the rewriting and translation of code, the conversion of code to text for the purposes of documentation and information retrieval, and the synthesis and general generation of code from text. Karampatsis et al. (2020) adds further applications of AI assistance such as code completion, API migration and code repair.
@@
Furthermore, Allamanis et al. (2018) presents [t]he naturalness hypothesis[:] Software is a form of human communication; software corpora have similar statistical properties to natural language corpora; and these properties can be exploited to build better software engineering tools.”. Given the intrinsic formats and formal characteristics of coding and modelling languages employed in SE, the second aspect of the naturalness hypothesis can be considered relatively straightforward. Nevertheless, the initial proposition reinforces the necessity for SE artefacts that are readily comprehensible. This assertion was previously made by  Fowler (2005) in a different form: “[A]ny fool can write code that a computer can understand, good programmers write code that humans can understand.”. Consequently, the application of AI in SE is not merely concerned with code generation; it also encompasses the enhancement of code through techniques such as refactorings, with the objective of optimising readability and maintainability.

### v2

# Ina K. Schieferdecker Technische Universität Berlin, Einsteinufer 25, 10587 Berlin, Germany

@@ Abstract
The effectiveness of model-driven software engineering (MDSE) has been successfully demonstrated in the context of complex software; however, it has not been widely adopted due to the requisite efforts associated with model development and maintenance, as well as the specific modelling competencies required for MDSE. Concurrently, artificial intelligence (AI) methods, particularly deep learning methods, have demonstrated considerable abilities when applied to the huge code bases accessible on open-source coding platforms. The so-called big code provides the basis for significant advances in empirical software engineering, as well as in the automation of coding processes and improvements in software quality with the use of AI. The objective of this paper is to facilitate a synthesis between these two significant domains of software engineering (SE), namely models and AI in SE. The paper provides an overview of the current state of AI-augmented software engineering and develops a corresponding taxonomy, AI4SE. In light of the aforementioned considerations, a vision of AI-assisted Big Models in SE is put forth, with the aim of capitalising on the advantages inherent to both approaches in the context of software development. Finally, the new paradigm of pair modelling in MDSE is proposed.
@@
Software Engineering, Model-Driven Engineering, Modelling, Domain Modelling, Artificial Intelligence, Big Models, AI4SE
@@ 1 Introduction
"Programming at the beginning of 2025 and at the end of 2025 are going to be completely different things."
Sam Altman, Febr. 7, 2025, The Age of AI – Panel Discussion, TU Berlin, Germany
@@
The advent of artificial intelligence (AI), and more specifically, generative AI, has precipitated a substantial transformation in the field of software engineering (SE), which is expected to undergo further significant changes in the near future (see also the quotation by Sam Altman, OpenAI). In order to better understand and comprehend these developments, this paper analysis the state of art in using AI for SE Software engineering is a field of informatics/computer science that addresses the development and analysis of systematic approaches for the design, development, verification & validation and maintenance of software and of software-based systems [^1] In this paper, the term ”software” will be used as a general reference to the collective term for computer programs and related SE artefacts.. It establishes principles and identifies optimal practices for the production and operation of software, with the aim of ensuring the reliability, security, scalability, maintainability and alignment of software with user, business and societal requirements. The objective of software engineering is the production of high-quality software that is cost-effective, delivered in a timely manner and can be readily adapted as requirements evolve.
@@
The term software engineering was coined in 1968 with the very first software crisis [82], but still software engineering practices and the resulting software quality have not kept pace with the quality levels required in critical domains or application contexts. In 2024, another dramatic story was added to the software horror show: The millions of enterprise outages caused worldwide by the CrowdStrike Falcon sensor update [24] were due to simple errors such as not validating the number of fields in an IPC template type, missing a runtime array bounds check, a logic error in the content validator, not testing template types enough, and not testing template instances within the content interpreter. This was unprofessional software development that ignored state of the art of SE.
@@
However, software outages are not the only indication that SE still has no full answers to the challenges posed by the increasing complexity of software-based systems and the diversity of requirements placed on them: [71] analysed that "[t]here exists a statistically significant medium sized difference between open and closed source projects: the former have a DD [(defect density)] that is 4 defects per KLoC lower than the latter. Java projects exhibit a significantly lower DD than C projects, 4.1 defects per KLoC on average: In general the [code] size appears to be negatively correlated to DD: the larger the project the lower the DD. In particular, large projects are 10 times less defective than medium ones.". According to [59], the defect density (DD) of software projects within the industrial sector is estimated to fall within the range of 1 to 25. This suggests that a software program comprising one million lines of code (MLoC) may contain between 1,000 and 25,000 defects. Given the lines of code in a piece of software, deriving estimates of the approximate number of defects in it is a relatively straightforward process. To illustrate, [74] compares typical software such as operating systems, browsers, and office suites: The Windows 10 operating system has approximately 80 MLoC, Ubuntu 50 MLoC, MacOS X 84 MLoC, Android 12 MLoC, iOS 12 MLoC, the browser Google Chrome 6.7 MLoC, Mozilla Firefox 21 MLoC, or the office suites Microsoft Office 2013 45 MLoC, Apache OpenOffice 19 MLoC, and LibreOffice 10 MLoC. Therefore, the challenges associated with the development of accurate and suitable software remain significant.
@@
[17] argued that the primary challenge in software development lies in the specification, design, and testing of complex systems, rather than in the labour for representation of the system or testing its fidelity. It emphasized the distinction between essential difficulties (inherent complexity) and accidental difficulties (extraneous challenges), stating that past advances, like high-level programming languages, have only reduced the latter. The paper suggested that addressing accidental difficulties would not lead to major improvements in software development, as essential difficulties remain fundamental. [36] critiqued the notion of "accidental difficulties", arguing that these so-called accidents are often the result of negligence or poor practices, not mere happenstance. This paper advocated for a disciplined, science-based, and model-driven approach to software development, similar to traditional engineering disciplines. Building on this, [68] highlighted the growing societal dependence on autonomous and intelligent software systems, suggesting that new approaches are needed to ensure not only traditional software quality facets (safety, security, etc.), but also to address socio-technical and socio-political implications, particularly in human-machine collaboration.
@@
In model-driven development (MDD [70]), also known as model-driven software engineering (MDSE [69]), models are original artefacts that are engineered with the intention of facilitating the top-down construction of complex software. Furthermore, models are utilised during runtime to enable the monitoring, verification & validation of software operations, as well as the optimisation of its performance [41, 34, 85]. The issues in using traditional (top-down) MDSE has been e.g. analysed in [35].
@@
As has been the case with the comprehensive incorporation of modelling in software engineering, AI techniques have been employed in software engineering activities from the outset, see for example [11]. In the context of significant advancements in AI-driven software engineering in 2021, Gartner’s hype cycle for emerging technologies [19] projected that AI-enhanced software engineering would reach the peak of inflated expectations within a five-to-ten-year timeframe, subsequently entering a phase of productivity stabilization. Since 2023, it is estimated that AI-augmented software engineering will reach the productivity plateau already in two to five years, while in previous iterations of the hype cycle, AI-augmented development and AI-assisted design were identified as emerging technologies on the rise. It was interesting to observe the evolving business expectations and the convergence of trends for AI-augmented design and development into a unified technological trend for AI-augmented engineering. Moreover, it is encouraging to observe that software engineering is receiving particular attention and prioritisation, which is indicative of the growing importance of this field.
Therefore, this paper focuses on the intersection of AI and SE, specifically MDSE, to improve software engineering practices. The paper aims to bridge the gap between the theoretical advantages of MDSE and its practical limitations by integrating AI techniques, and introduces a novel taxonomy, AI4SE, to classify AI applications within the software engineering domain, highlight open research questions, and direct future research. The paper uses a combination of literature review, analysis of existing software engineering practices and AI applications, and development of a novel taxonomy to achieve its objectives. The main contributions of the paper include:
  • The new AI in Software Engineering Taxonomy AI4SE. This taxonomy categorises AI applications based on the purpose of using AI, the target of AI within software engineering (e.g. requirements, design, coding, testing, operations, process), the type of AI used, and the level of autonomy achieved by the AI.
  • The novel Modelling Hypothesis, which posits that models have statistical properties that can be used to develop more effective software engineering tools.
  • The concept of Big Models, analogous to big code, as a basis for advanced empirical MDSE. The paper suggests that these models can be generated top-down, or extracted bottom-up from code, and that AI methods can be applied to these big models.
  • The new paradigm of pair modelling in MDSE, where a software engineer and an AI tool/agent collaborate in software development, similar to pair programming. The aim is to leverage the strengths of both human and AI for MDSE to improve software quality.
The AI4SE taxonomy is validated by categorising highly cited and recent publications in AI-assisted software engineering. This classification shows the various ways AI is being applied across different activities in SE, such as requirements engineering, architecture design, coding, testing, and deployment.
@@
In section 2, the paper addresses the role of models in software engineering. The evolution of big code on coding platforms is discussed in section 3 and the utilisation of AI for coding and software engineering in section 4. The novel taxonomy of AI for SE (AI4SE) to facilitate a deeper understanding of this evolving field is presented also in this section. Section 5 examines the practice of contributing software models to coding platforms, which is increasingly being done together with software code. In line with the increasing prevalence of large-scale models, it examines recent developments in AI-enhanced modelling and model utilisation within the domain of software engineering and presents the Big Model concept and the pair modelling paradigm. In section 6, related work is reviewed. A summary concludes the paper.
@@ 2 Models in Software Engineering
According to the IEEE Software Engineering Body of Knowledge [73] (see Figure 1), the software lifecycle is comprised of distinct phases and activities that can be described as knowledge areas, including requirements, architecture, design, construction, testing and maintenance. There are also knowledge areas that deal with the fundamentals of computer science, mathematics and engineering, as well as cross-cutting activities pertaining to software quality, security, configuration management and engineering management. Furthermore, there are cross-cutting knowledge areas that encompass the processes, operations, professional practice and economics of software engineering, along with a distinct area for software engineering models and methods. This knowledge area encompasses the processes and methodologies associated with modelling, the various types of models, including those pertaining to information, behaviour and structure, and the analysis of models, see also [69]. Overall, SWEBOK addresses dedicated model-based or model-driven approaches in each knowledge area, such as model-based requirements specification, model-based architecture, model-driven design, lifecycle models, executable models, model-based testing, etc.
Figure 1: Software Engineering knowledge areas according to SWEBOK [73]
This is because models in SE are the key to dealing with the complexity of software: Models provide the essential abstractions to capture requirements, support design decisions, or to offer comprehensive overviews of software structures and behaviours. They are essential tools in the engineering process for constructing and maintaining software. They are also vital for configuration, monitoring and other runtime support during software operation and management. The understanding of Model-Driven Software Engineering (MDSE) has evolved in recent years as the capabilities of computing, modelling and the available software and data have evolved:
@@
    In addition to explicit model development and top-down model use, models can also be extracted from software executions, from software monitorings or from code artefacts with the objective of avoiding the costs associated with model development on the one hand and leveraging the benefits of MDSE on the other hand [7].
@@
    The combination of model development of model artefacts with model extraction from code artefacts has resulted in an hybrid top-down/bottom-up MDSE approaches [76] benefiting from both directions of MDSE as shown in [83, 85] or [37].
@@
    Furthermore, as the abstractions inherent in programming became more intricate, low-level information and structure models have become indispensable to conventional programming. This phenomenon is exemplified by the utilisation of data structures that are frequently defined in JSON, XML schemata, or SQL table specifications, to name but a few.
These practical outcomes of model-driven, model-based and/or model-oriented software developments established a foundation for the utilisation of AI in software engineering, as discussed in Section 5. Before that, however, section 3 will take a closer look at the current state of big code on open source platforms, and section 4 will take a closer look at the state of applying AI to big code.
@@ 3 Big Code in Software Engineering
The advent of software coding platforms can be traced back to the 1990s, although it was not until the early 2000s that they truly gained prominence, with the launch of GitHub in April 2008 representing a significant milestone in this regard. At the present time, GitHub is the most widely utilised software coding platform. As stated by [15], GitHub hosts over 100 million projects and 40 million users. It became not only a significant platform for software engineering collaboration, but also a prominent reference for open-source software mining [45]. The study also demonstrated that a considerable number of GitHub repositories are not directly related to software development. This is because GitHub is not solely utilised for coding collaborations; it is also employed for collaborations on websites, editing books or other publications, and is even used as a storage platform. A later manual analysis by the same authors [46] revealed that over a third of the repositories on GitHub were not software development repositories, which degraded now to roughly 20 percent according to [4] (see figure 2 upper left).
Nevertheless, the software engineering community can further improve its understanding of software engineering processes, methods and tools by gaining insights through the mining of coding platforms using tools such as Kibble [4]. In a Feb. 2025 Kibble snapshot analysing lines of code from all GitHub code repositories, about 70% is actual code, while the other 30% is comments and empty lines. Java is the most common coding language, followed by XML and JSON for data and data structures (see figure 2 below left). A GitHub search counting files in GitHub repositories shows the ratio of the number of files for coding files to model files for software designs, interface models and data models (see figure 2 top right). The split between design models, interface models and data models is shown in figure 2 below right, which shows that about 1% of the files on GitHub contain models or partial model information.
Figure 2: Repositories breakdown by Kibble and GitHub Search taken 4-6 February 2025
All of these figures are only rough indicators: Lines of code and number of files are crude measures of software. File extensions do not accurately indicate file types. Coding and modelling tools are numerous and sometimes have proprietary file extensions. Overall, a GitHub search for more specific file types such as UML works better than for more general file types such as XML or JSON, but can still lead to false positives.
Nevertheless, the sheer numbers show that modelling is being used in software development. A more detailed approach to modelling on GitHub, including image analysis, is given in [72], but this is not applicable to GitHub as a whole due to API limitations. Despite this, the numbers show that modelling is being actively used. In fact, modelling is mainly used for data/interface structures, due to the simple fact that data and interfaces do not work without at least a well-defined structure. In contrast, design and behaviour models for complete software (packages) are only used for a small fraction.
This is very much in line with the analysis of [34] over a decade ago: Code generators may produce a large amount of code, but not all of it is necessarily useful or would have been written manually. The effort required to develop, maintain and adapt the generators reduces the productivity gains from modelling. When manual code changes are required, they make round-trip engineering difficult. Generated code must be checked by developers to ensure correctness, limiting the benefits of automation. Poor software design models can lead to inefficient development. As a result, developers often prefer generic solutions with hand-written, reusable code to customised, generated solutions. As programming languages and frameworks become more expressive, the need for model-driven code generation diminishes. In this sense, GenAI further empowers the software developer. But does this mean that MDSE will disappear, or will the ability to (re)generate models from code and/or keep them in sync with software requirements evolution and manual code changes result in a boost for MDSE (see section 5)?
@@
Back to big code: Due to the large number of software repositories with roughly 500 MB code each, the concept of "Big Code", analogous to the notion of "Big Data", was introduced in reference to the vast quantities of software code that have accumulated over time [58, 61, 84]. This code can be reused, extended and analysed to gain deeper insights into software engineering. It can be used to train ML models and enabled the application of AI for SE in large scale [53].
4 Artificial Intelligence in Software Engineering
@@
The use cases for AI in SE are numerous and cover a wide spectrum as discussed already back in 1988 [11] and since then [2]. For example, [31] proposes the AI in SE Application Levels (AI-SEAL) taxonomy, which differentiates between the point of applying AI to the software engineering process, the software product or at runtime, the levels of automation, and the types of AI along the five tribes differentiation by [27]. It is noteworthy that the majority of the papers analysed in [31] employ AI in the software engineering process. However, despite the numerous facets of the software engineering process (Figure 1), this study does not provide further elaboration. Nevertheless, other publications offer more detailed discussions of AI in SE, including [10] or [62]. Notwithstanding the aforementioned considerations, a taxonomy for the role of AI in software engineering has yet to emerge.
@@
In light of the recent advancements in machine learning (ML) that have made significant breakthroughs possible, this paper focuses mainly on the latest developments in the application of ML to SE. For ML to be effective, it is essential to utilise the appropriate structures inherent to SE artefacts, which has been a topic of considerable debate: [3] presents an overview of the various ways in which source code can be represented, including representational models of tokens, token contexts, program dependency graphs, API calls, abstract syntax trees, object usage, and others. These representational models are used for AI assistance in SE, including the creation of recommender systems, the inference of coding conventions, the detection of anomalies and defects, the analysis of code, the rewriting and translation of code, the conversion of code to text for the purposes of documentation and information retrieval, and the synthesis and general generation of code from text. [47] adds further applications of AI assistance such as code completion, API migration and code repair.
@@
Furthermore, [3] presents "[t]he naturalness hypothesis[:] Software is a form of human communication; software corpora have similar statistical properties to natural language corpora; and these properties can be exploited to build better software engineering tools.". Given the intrinsic formats and formal characteristics of coding and modelling languages employed in SE, the second aspect of the naturalness hypothesis can be considered relatively straightforward. Nevertheless, the initial proposition reinforces the necessity for SE artefacts that are readily comprehensible. This assertion was previously made by [33] in a different form: "[A]ny fool can write code that a computer can understand, good programmers write code that humans can understand.". Consequently, the application of AI in SE is not merely concerned with code generation or code repair; it also encompasses the enhancement of code through techniques such as refactorings, with the objective of optimising readability and maintainability.